# Correlated Growth of Causal Networks

### v1

@@ Abstract
The study of causal structure in complex systems has gained increasing attention, with many recent studies exploring causal networks that capture cause-effect relationships across diverse fields. Despite increasing empirical evidence linking causal structures to network topological correlations, the mechanisms underlying the emergence of these correlations in causal networks remain poorly understood. In this Letter, we propose a general growth framework for causal networks, incorporating two key types of correlations: causal and dynamic. We analytically solve our model for degree correlation and validate the theoretical predictions against empirical data from four large-scale innovation networks. Our theory not only sheds light on the origins of topological correlations but also provides a general framework for understanding correlated growth across causal systems.
@@
A causal network, from a physics perspective, is inherently rigid due to the immutability of past events. Nodes represent events, and directed edges denote fixed causal relationships. Unlike general networks, causal networks cannot be rewired or retroactively modified without violating temporal and logical consistency. This rigidity can be represented by directed acyclic graphs (DAGs), reflecting the unidirectional arrow of time. Consequently, causal networks serve as static records of chronological and causal relationships, distinguishing them from dynamic, modifiable general networks [21, 22, 23]. For example, while general networks like social media follower-followee relationships can evolve dynamically via rewiring or deletion, causal networks provide static records of interactions such as “follow” or “unfollow,” preserving the historical flow of actions. In causal networks, such dynamics are simplified through an event creation mechanism, leveraging the growth network approach [24, 25, 13, 26, 1], enabling broader applicability.
@@
Causal networks are characterized by various correlations that quantify dependencies between entities. Of particular importance are topological correlations, which reveal the non-random organization of the network structure [27, 28, 29, 30]. These correlations provide critical insights into how causal relationships shape the network architecture. Empirical evidence increasingly suggests that topological correlations are crucial to the development of causal networks [31, 32, 1]. For instance, topological correlations in citation networks are believed to govern the inheritance patterns of scientific impacts [33]. Similarly, in ecological event networks, topological correlations that reflect the causality between different events play a key role in shaping microbial community structures [34]. While there has been substantial progress in quantifying topological correlations for general networks [35, 36, 29, 37], relatively few studies have attempted to model network correlations explicitly [38, 39, 40]. Moreover, these models primarily focus on non-causal networks. The fundamental distinction in the causal network implies that traditional network correlation models, which rely on dynamic rewiring mechanisms [38, 39, 40], cannot fully capture the unique correlation patterns that arise within causal networks.
@@
In this letter, we propose a general correlated growth framework for casual networks, incorporating two key mechanisms: 1) Causal correlation, where causal relationships between individuals and their successors are transmitted across generations; and 2) Dynamic correlation, which captures the long-range temporal correlations during the growth. We analytically solve the proposed model for the degree correlation, a widely used metric for quantifying topological correlations. By applying our theory to large-scale innovation networks across four disciplines, we show that our theoretical predictions agree quantitatively with empirical observations. As a unified approach for modeling correlated growth in causal networks, our theory offers the potential to impact a diverse range of fields.
@@
Theoretical Framework. Consider a causal network model that starts with $n_{0}$ events at time $t=0$. As the network evolves, each newly added event $i$ connects to preexisting events introduced at earlier times, yielding a directed acyclic graph (DAG) that ensures causality. Additionally, each event $i$ is assigned a state $\mathbf{s}_{i}$, which may include both discrete and continuous attributes. For instance, in a social network, $\mathbf{s}_{i}$ could encode the type of interaction (discrete) alongside its intensity or duration (continuous). To capture a broad range of scenarios, we separate $\mathbf{s}_{i}$ into two subspaces: an observed component $\mathbf{s}_{\mathrm{obs}}$, reflecting publicly accessible properties, and a latent component $\mathbf{s}_{\mathrm{lat}}$, encoding hidden variables that influence network evolution. One key observed trait is $m_{i}\in\mathbf{s}_{\mathrm{obs}}$, the number of existing events that a newly created event $i$ connects to.
@@
Once an event with state $\mathbf{s}$ is created at time $t$, its causal relationships can continue to grow. We capture this individual-level growth via a Green’s function $G_{\mathbf{s}}(k,\tau|k^{\prime},\tau^{\prime})$, which gives the probability that an event with $k^{\prime}$ connections at age $\tau^{\prime}$ grows to $k$ connections at a later age $\tau$. Here, $\tau^{\prime}\mathrel{\mathop{:}}=t^{\prime}-t_{i}$ and $\tau\mathrel{\mathop{:}}=t-t_{i}$, measured from the event’s creation time $t_{i}$. As an example, a Poisson process with rate $\lambda$ yields $G_{\lambda}(k,\tau|k^{\prime},\tau^{\prime})=\frac{\left[\lambda(\tau-\tau^{\prime})\right]^{k-k^{\prime}}e^{-\lambda(\tau-\tau^{\prime})}}{(k-k^{\prime})!}$. However, $G$ remains a general input to our framework, accommodating a wide range of growth processes. The mean accumulative connection, $\overline{k}(\mathbf{s},\tau)\mathrel{\mathop{:}}=\sum_{k}kG_{\mathbf{s}}(k,\tau|0,0)$, then tracks how an event’s expected degree evolves over its lifetime.
@@
Global network growth arises from the collective growth of individual events. To describe this process, we focus on $n(\mathbf{s},t)$, the rate at which new events with state $\mathbf{s}$ appear. Directed edges from newly introduced vertices to existing ones encode causal correlations; for example, an “unfollow” action can only occur after a corresponding “follow.” To capture these correlations, we employ a mean-field causal kernel $K(\mathbf{s},t|\mathbf{s}^{\prime},t^{\prime})$, which gives the probability that a vertex created at time $t$ has state $\mathbf{s}$ given that it links to a vertex with state $\mathbf{s}^{\prime}$ created at an earlier time $t^{\prime}<t$. The event-creation rate then satisfies
@@
$\frac{1}{\langle k\rangle}\int_{0}^{t}\mathrm{d}t^{\prime}\int_{0}^{\infty}\mathrm{d}\mathbf{s}^{\prime}\;n(\mathbf{s}^{\prime},t^{\prime})K(\mathbf{s},t|\mathbf{s}^{\prime},t^{\prime})\partial_{t}\overline{k}(\mathbf{s}^{\prime},t-t^{\prime}),$ (1)
where $\langle k\rangle$ is the average degree. For constant $m_{i}=m$, we have $\langle k\rangle=m$.
To solve this equation, we focus primarily on its stationary solution. In this regime, $n(\mathbf{s},t)$ factorizes asymptotically as $n(t)\,\psi(\mathbf{s})$, where $\psi(\mathbf{s})$ is the stationary state distribution, and $n(t)\equiv\int n(\mathbf{s},t)\,\mathrm{d}\mathbf{s}\;=\;\frac{\mathrm{d}N(t)}{\mathrm{d}t}$ is the total growth rate of the network size. One can show that the existence of such a stationary solution requires the growth process to exhibit time-translation invariance (TTI), i.e. $K(\mathbf{s},t|\mathbf{s}^{\prime},t^{\prime})=K(\mathbf{s}|\mathbf{s}^{\prime},t-t^{\prime})$ and $n(t)/n(t^{\prime})\sim f(t-t^{\prime})$, so that only time differences matter. Under these conditions, we predict an exponential growth
@@
where $r$ is the network growth rate. Substituting this form into Eq. (Correlated Growth of Causal Networks) and taking $t\to\infty$ yields an analytic solution for the stationary distribution
$\psi(\mathbf{s})=\frac{1}{\langle k\rangle}\int\mathrm{d}\tau\int\mathrm{d}\mathbf{s}^{\prime}K(\mathbf{s}|\mathbf{s}^{\prime},\tau)\partial_{\tau}{\overline{k}(\mathbf{s}^{\prime},\tau)}e^{-r\tau}\psi(\mathbf{s}^{\prime}),$ (3)
@@
If Equation (3) admits no solution, stationarity breaks down. This can happen when $\mathbf{s}$ is unbounded and $K(\mathbf{s}|\mathbf{s}^{\prime},\tau)$ has sufficiently broad tails, causing events to attain arbitrarily large values of $\mathbf{s}$ and thereby driving super-exponential growth of $n(t)$, violating TTI. This scenario is reminiscent of a “winner-takes-all” phase, analogous to Bose-Einstein condensation [25]. Since this behavior does not arise in the empirical data we examine later, we defer its exploration to future work and focus on the stationary case.
Once $\psi(\mathbf{s})$ and $r$ are obtained from Eq. (3), we can analytically compute related network characteristics. For a given observable $Q(\mathbf{s})$, we project it onto the observed subspace via
@@
$\langle Q(\mathbf{s}_{\text{obs}})\rangle_{\text{lat}}\mathrel{\mathop{:}}=\int d\mathbf{s}_{\text{lat}}\;Q(\mathbf{s}_{\text{obs}},\mathbf{s}_{\text{lat}})\psi_{\text{lat}}(\mathbf{s}_{\text{lat}}),$ (4)
where $\psi_{\mathrm{lat}}(\mathbf{s}_{\mathrm{lat}})\equiv\int\mathrm{d}\mathbf{s}_{\mathrm{obs}}\,\psi(\mathbf{s}_{\mathrm{obs}},\mathbf{s}_{\mathrm{lat}})$.
A key network characteristic is the degree distribution $P(k|\mathbf{s})$, which measures the probability that a vertex with state $\mathbf{s}$ has degree $k$. Because $n(t)\sim e^{rt}$, the age distribution of vertices follows $P(\tau)=r\,e^{-r\tau}$. Integrating over $\tau$ and $\mathbf{s}$ then gives
$P(k|\mathbf{s})=\int_{0}^{\infty}\mathrm{d}\tau\;G_{\mathbf{s}}(k,\tau|0,0)re^{-r\tau},$ (5)
where $G_{\mathbf{s}}(k,\tau|0,0)$ is the probability that a vertex with state $\mathbf{s}$ and zero initial links grows to degree $k$ over age $\tau$.
To quantify degree correlations, we examine the joint distribution $P(k,\mathbf{s};k^{\prime},\mathbf{s}^{\prime})$, the probability of finding an edge from a node of degree $k^{\prime}$ and state $\mathbf{s}^{\prime}$ to a node of degree $k$ and state $\mathbf{s}$. We find
@@
$P(k,\mathbf{s};k^{\prime},\mathbf{s}^{\prime})=\frac{1}{\langle k\rangle}$ $\int_{0}^{\infty}\mathrm{d}\tau^{\prime}\mathrm{d}\tau\;\psi(\mathbf{s}^{\prime})re^{-r\tau^{\prime}}P(k^{\prime},\tau|\mathbf{s}^{\prime},\tau^{\prime})$
×$\times K(\mathbf{s}|\mathbf{s}^{\prime},\tau^{\prime}-\tau)G_{\mathbf{s}}(k,\tau|0,0),$ (6)
where $P(k^{\prime},\tau|\mathbf{s}^{\prime},\tau^{\prime})$ is the probability that a vertex born at $t-\tau^{\prime}$ has degree $k^{\prime}$ at age $\tau$, fully determined by the Green’s function $G$. In the absence of both causal and dynamical correlations (i.e. $K(\mathbf{s}|\mathbf{s}^{\prime},\tau^{\prime}-\tau)=\psi(\mathbf{s})$ and no growth dependence on state), one can show that $P(k,k^{\prime})$ factorizes.
@@
Empirical Validation. To validate our theory against empirical data, we applied it to the scientific innovation system, which can be understood as a causal network, where the creation of innovations causally depends on existing knowledge. We used the Web of Science (WOS) dataset spanning 1970–2014, comprising 44M publications and 800M citations. For simplicity, the causal network considered here is the citation network, focusing on events of paper publication with a single type of causal relationship $s_{\text{obs}}=$ “cite”, which links new knowledge to its intellectual predecessors. Using the WOS dataset, we extracted four citation networks across different disciplines, including biology, chemistry, mathematics, and physics, comprising 13M, 9M, 3M, and 7M nodes, respectively .
@@
Figure 1: Using data from the Web of Science, a measures $n(t)$ versus $t$ for the citation graphs across biology, chemistry, mathematics, and physics. The slopes of the solid lines represent the theoretical predictions of the growth rate $r$. b shows the empirical data (scatter points) and model predictions (solid lines) for the probability distribution of states. c presents the empirical data (scatter points) and model predictions (solid lines) for the degree distribution across the four subjects. d provides an example of the normalized causal correlation kernel $\overline{\mathbf{s}}(\mathbf{s}^{\prime},\tau)^{-1}K(\mathbf{s}|\mathbf{s}^{\prime},\tau)=\kappa({\mathbf{s}}/{\overline{\mathbf{s}}}(\mathbf{s}^{\prime},\tau))$ for physics. The corresponding $\kappa({\mathbf{s}}/{\overline{\mathbf{s}}}(\mathbf{s}^{\prime},\tau))$ for other subjects is shown in the SM.
Existing studies suggest that the growth process at the individual level follows a reinforced Poisson process (RPP), incorporating both preferential attachment and aging [41, 12]. Specifically, the growth rate is given by $\Lambda{(k,\tau)}=\lambda\phi(\tau)(k+k_{0})$, where $\lambda$ represents fitness. The aging function, typically modeled as a log-normal distribution $\phi(t)=\frac{1}{\sqrt{2\pi}\sigma t}\exp\left[-\frac{(\ln{t}-\mu)^{2}}{2\sigma^{2}}\right]$, captures the long-tail dynamical correlation. The corresponding Green’s function is given by the negative binomial distribution, with the average growth
${\overline{k}}(\lambda,t)=k_{0}\left(\exp\left(\lambda\int_{0}^{t}\phi(t^{\prime})\mathrm{d}t^{\prime}\right)-1\right).$ (7)
While Eq. (7) successfully captures individual citation growth, it has several drawbacks: (1) it does not account for causal correlations, and (2) it assigns the parameters $\lambda$, $k_{0}$, $\mu$, and $\sigma$ to each paper, leading to a large number of free parameters that are experimentally determined. In particular, the number of parameters scales with the number of vertices, $O(N)$. This proliferation of parameters diminishes the model’s predictive and explanatory power by introducing excessive degrees of freedom. Such limitations are well-documented in fitness models [24, 25].
@@
Figure 2: normalized joint degree distribution. This figure shows the empirical data and modeling predictions of $k_{nn}(k_{0})$ across biology a&b, chemistry c&d, math e&f, and physics g&h.
@@
Our theory addresses the above drawbacks by treating individual parameters as latent states, $\mathbf{s}_{\text{lat}}$. This approach allows our theory (3) to predict the distribution of these latent states, thereby avoiding the large parameter space. To solve Eq. (3), it is necessary to determine the causal correlation kernel, $K$. For simplicity, we focus solely on the fitness parameter $\mathbf{s}_{\text{lat}}=\lambda$, keeping all other parameters fixed globally. Figure 1d plots a rescaled $K(\lambda|\lambda^{\prime},\tau)$, which is well approximated by a stretched exponential. Furthermore, it collapses into master curves after scaling by its mean $\overline{\lambda}$, as
$K(\lambda|\lambda^{\prime},\tau)=\overline{\lambda}(\lambda^{\prime},\tau)\kappa\left({\lambda}/{\overline{\lambda}}(\lambda^{\prime},\tau)\right)$ (8)
Here, the mean fitness $\overline{\lambda}(\lambda^{\prime},\tau)$ is determined by Eq. 7, as shown in our previous work [42], and $\kappa(x)\sim x^{\beta-1}e^{-(x/b)^{\beta}}$ is the stretched exponential function . Thus, our theory significantly reduces the parameter space, from being dependent on individual papers to a small set of global parameters, thereby decreasing complexity from $O(N)$ to $O(1)$.
@@
Substituting Eqs.(7) and (8) into Eq.(3), our theory predicts the network growth rates $r\approx 0.037$, $0.03$, $0.04$, and $0.028$ for biology, chemistry, mathematics, and physics, respectively, which are in excellent agreement with the empirical measurements (Fig. 1a). Figure 1b compares the predicted fitness distribution $\psi(\lambda)$ with the empirical data, showing remarkable consistency. Notably, the empirical $\psi(\lambda)$ is obtained by fitting the PRR model to individual papers, whereas our theory provides a direct prediction without requiring individual fitness values. Additionally, using Eqs. (4)–(5), the degree distribution $P(k)=\langle P(k|\lambda)\rangle_{\lambda}$ is computed analytically, again aligning perfectly with experimental results (Fig. 1c).
It is well-known that the citation graph is assortative [35]. To validate the model’s prediction of topical correlation, we calculate the joint degree distribution

### v2

@@ Abstract
The study of causal structure in complex systems has gained increasing attention, with many recent studies exploring causal networks that capture cause-effect relationships across diverse fields. Despite increasing empirical evidence linking causal structures to network topological correlations, the mechanisms underlying the emergence of these correlations in causal networks remain poorly understood. In this work, we propose a general growth framework for causal networks, incorporating two key types of correlations: causal and dynamic. We analytically demonstrate that degree correlations emerge as a consequence of marginal dependencies on these correlations. Our theoretical predictions align quantitatively with empirical data from four large-scale innovation networks. Our theory not only sheds light on the origins of topological correlations but also provides a general framework for understanding correlated growth across causal systems.
@@
A causal network, from a physics perspective, is inherently rigid due to the immutability of past events. Vertices represent events, and directed edges denote fixed causal relationships. Unlike general networks, causal networks cannot be rewired or retroactively modified without violating temporal and causal consistency. This rigidity can be represented by directed acyclic graphs (DAGs), reflecting the unidirectional arrow of time. Consequently, causal networks serve as static records of chronological and causal relationships, distinguishing them from dynamic, modifiable general networks [21, 22, 23]. For example, while general networks like social media follower-followee relationships can evolve dynamically via rewiring or deletion, causal networks provide static records of interactions such as “follow” or “unfollow,” preserving the historical flow of actions. In causal networks, such dynamics are simplified through an event creation mechanism, leveraging the growth network approach [24, 25, 13, 26, 1], enabling broader applicability.
@@
Causal networks are characterized by various correlations that quantify the dependencies between entities. Of particular importance are topological correlations, which reveal the non-random organization of the network structure [27, 28, 29, 30]. These correlations provide critical insights into how causal relationships shape the network architecture. Empirical evidence increasingly suggests that topological correlations are crucial for the development of causal networks [31, 32, 1]. For instance, topological correlations in citation networks are believed to govern the inheritance patterns of scientific impacts [33]. Similarly, in ecological event networks, topological correlations that reflect the causality between different events play a key role in shaping microbial community structures [34]. While there has been substantial progress in quantifying topological correlations for general networks [35, 36, 29, 37], relatively few studies have attempted to model network correlations explicitly [38, 39, 40]. Moreover, these models primarily focus on non-causal networks. The fundamental distinction in the causal network implies that traditional network correlation models, particularly those relying on dynamic rewiring mechanisms [21, 22, 23], cannot fully capture the unique correlation patterns that arise within causal networks.
@@
In this letter, we propose a general correlated growth framework for casual networks, incorporating two key mechanisms: 1) Dynamic correlation, which captures the long-range temporal correlations during the individual degree growth; 2) Causal correlation, where causal relationships between individuals and their successors are transmitted across generations. By analytically solving the joint degree distribution, we explicitly demonstrate that degree correlations emerge from marginal dependencies on both dynamic and causal correlations. Applying our theory to large-scale innovation networks across four disciplines, we find strong quantitative agreement between our theoretical predictions and empirical observations. As a unified approach to modeling correlated growth in causal networks, our framework has the potential to impact a wide range of fields.
@@
Theoretical Framework. Consider a causal network model that starts with $n_{0}$ events at time $t=0$. As the network evolves, each newly added event $i$ connects to $m_{i}$ preexisting events introduced at earlier times, yielding a DAG that ensures causality. Additionally, each event $i$ is assigned a state $\mathbf{s}_{i}$, which may include both discrete and continuous attributes. For instance, in a social network, $\mathbf{s}_{i}$ could encode the type of interaction (discrete) alongside its intensity or duration (continuous). To capture a wide range of scenarios, we divide $\mathbf{s}_{i}$ into two subspaces: an observed component $\mathbf{s}_{\mathrm{obs}}$, reflecting publicly accessible properties, and a latent component $\mathbf{s}_{\mathrm{lat}}$, encoding hidden variables that influence network evolution.
@@
Once an event with state $\mathbf{s}$ is created at time $t$, its causal relationships can continue to grow. We capture this individual-level growth via a Green’s function $G_{\mathbf{s}}(k,\tau|k^{\prime},\tau^{\prime})$, which gives the probability that an event with $k^{\prime}$ connections at age $\tau^{\prime}$ grows to $k$ connections at a later age $\tau$. Here, $\tau^{\prime}\mathrel{\mathop{:}}=t^{\prime}-t_{i}$ and $\tau\mathrel{\mathop{:}}=t-t_{i}$, measured from the event’s creation time $t_{i}$. The corresponding mean accumulative connection, $\overline{k}(\mathbf{s},\tau)\mathrel{\mathop{:}}=\sum_{k}kG_{\mathbf{s}}(k,\tau|0,0)$, then tracks how an event’s expected connections evolve over its lifetime. As an example, a Poisson process with rate $\lambda$ yields $G_{\lambda}(k,\tau|k^{\prime},\tau^{\prime})=\frac{\left[\lambda(\tau-\tau^{\prime})\right]^{k-k^{\prime}}e^{-\lambda(\tau-\tau^{\prime})}}{(k-k^{\prime})!}$, and $\overline{k}(\mathbf{s},\tau)=\lambda\tau$. However, for a process with temporal correlation, the corresponding Green’s function $G$ may exhibit a fat tail rather than an exponential decay. In our framework, $G$ remains a general input to accommodate different growth processes.
@@
Global network growth arises from the collective growth of individual events. To describe this process, we focus on $n(\mathbf{s},t)$, the rate at which new events with state $\mathbf{s}$ appear. The directed edges from newly introduced vertices to existing ones encode causal correlations; for example, an “unfollow” action can only occur after a corresponding “follow.” To capture these correlations, we employ a mean-field causal kernel $K(\mathbf{s},t|\mathbf{s}^{\prime},t^{\prime})$, which gives the probability that a vertex created at time $t$ has state $\mathbf{s}$ given that it links to a vertex with state $\mathbf{s}^{\prime}$ created at an earlier time $t^{\prime}<t$. The event-creation rate then satisfies
@@
$\frac{1}{\langle k\rangle}\int_{0}^{t}\mathrm{d}t^{\prime}\int\mathrm{d}\mathbf{s}^{\prime}\;n(\mathbf{s}^{\prime},t^{\prime})K(\mathbf{s},t|\mathbf{s}^{\prime},t^{\prime})\partial_{t}\overline{k}(\mathbf{s}^{\prime},t-t^{\prime}),$ (1)
where $\langle k\rangle$ is the average degree. For simplicity, in the discussion below, we consider the case of a fixed outdegree, $m_{i}=m$, so that $\langle k\rangle=m$. In more general cases, $m_{i}$ can be incorporated into the state $\mathbf{s}_{i}$.
To solve Eq. (Correlated Growth of Causal Networks), we focus primarily on its stationary solution. In this regime, $n(\mathbf{s},t)$ factorizes asymptotically as $n(t)\psi(\mathbf{s})$, where $\psi(\mathbf{s})$ is the stationary state distribution, and $n(t)\equiv\int n(\mathbf{s},t)\,\mathrm{d}\mathbf{s}\;=\;\frac{\mathrm{d}N(t)}{\mathrm{d}t}$ is the total growth rate of the network size. One can show that the existence of such a stationary solution requires the growth process to exhibit time-translation invariance (TTI), i.e. $K(\mathbf{s},t|\mathbf{s}^{\prime},t^{\prime})=K(\mathbf{s}|\mathbf{s}^{\prime},t-t^{\prime})$ and $n(t)/n(t^{\prime})\sim f(t-t^{\prime})$, so that only time differences matter (see SM Section 1.2). Under these conditions, we predict an exponential growth
@@
where $r$ is the network growth rate. The analysis above also implies the asymptotic state-age distribution,
[…]
factorizes into the product of the state and age distributions.
Substituting Eq. (2) into Eq. (Correlated Growth of Causal Networks) and taking $t\to\infty$ yields leads the stationary master equation
$\psi(\mathbf{s})=\frac{1}{\langle k\rangle}\int_{0}^{\infty}\mathrm{d}\tau\int\mathrm{d}\mathbf{s}^{\prime}K(\mathbf{s}|\mathbf{s}^{\prime},\tau)\partial_{\tau}{\overline{k}(\mathbf{s}^{\prime},\tau)}e^{-r\tau}\psi(\mathbf{s}^{\prime}),$ (4)
@@
Network Characteristics. Once $\psi(\mathbf{s})$ and $r$ are obtained from Eq. (4), the state-age distribution $\Psi(\mathbf{s},\tau)$ can be determined using Eq. (3), enabling an analytical computation of various network characteristics. We now outline the computation of the degree distribution $P(k)$, which represents the probability of finding a vertex with in-degree $k$, and the joint degree distribution $P(k,k^{\prime})$, which gives the probability of finding an edge from a vertex with in-degree $k$ to a vertex with in-degree $k^{\prime}$. The latter is a fundamental measure that characterizes network correlations.
A key quantity in our framework for understanding these network characteristics is the joint state-age distribution,
@@
$\mathbf{\Psi}(\mathbf{s},\tau;\mathbf{s}^{\prime},\tau^{\prime})\mathrel{\mathop{:}}=\frac{1}{\langle k\rangle}K(\mathbf{s}|\mathbf{s}^{\prime},\tau^{\prime}-\tau)\partial_{\tau^{\prime}}{\overline{k}(\mathbf{s}^{\prime},\tau^{\prime}-\tau)}\Psi(\mathbf{s}^{\prime},\tau^{\prime}),$
which quantifies the probability of finding an edge from a vertex with state $\mathbf{s}$ and age $\tau$ to a vertex with state $\mathbf{s}^{\prime}$ and age $\tau^{\prime}$ ($\tau^{\prime}>\tau$). Using Eq. (4), one obtains the marginal probabilities at either of these two vertices: $\int\mathbf{\Psi}(\mathbf{s},\tau;\mathbf{s}^{\prime},\tau^{\prime})\mathrm{d}\mathbf{s}^{\prime}\mathrm{d}\tau^{\prime}=\Psi(\mathbf{s},\tau)$ and $\int\mathbf{\Psi}(\mathbf{s},\tau;\mathbf{s}^{\prime},\tau^{\prime})\mathrm{d}\mathbf{s}\mathrm{d}\tau=\overline{k}(\mathbf{s}^{\prime},\tau^{\prime})\Psi(\mathbf{s}^{\prime},\tau^{\prime})/\langle k\rangle.$
Since $G_{\mathbf{s}}(k,\tau|0,0)$ gives the probability that a vertex with state $\mathbf{s}$ and age $\tau$ attains in-degree $k$, the degree distribution is given by $P(k)=\int\mathrm{d}\mathbf{s}\int_{0}^{\infty}\mathrm{d}\tau\;G_{\mathbf{s}}(k,\tau|0,0)\Psi(\mathbf{s},\tau),$ which can be interpreted as the averaged Green’s function,
$P(k)=\left\langle G_{\mathbf{s}}(k,\tau|0,0)\right\rangle,$ (5)
where the average $\left\langle\ldots\right\rangle\mathrel{\mathop{:}}=\int\ldots\mathbf{\Psi}(\mathbf{s},\tau;\mathbf{s}^{\prime},\tau^{\prime})\mathrm{d}\mathbf{s}\mathrm{d}\tau\mathrm{d}\mathbf{s}^{\prime}\mathrm{d}\tau^{\prime}$ is taken over the joint state-age distribution. Similarly, one finds (see SM Section 1.3)
@@
$P(k,k^{\prime})=\left\langle\frac{k^{\prime}}{\overline{k}(\mathbf{s}^{\prime},\tau^{\prime})}G_{\mathbf{s}^{\prime}}(k^{\prime},\tau^{\prime}|0,0)G_{\mathbf{s}}(k,\tau|0,0)\right\rangle,$ (6)
where $k^{\prime}$ represents the number of edges connecting to the vertex, and $\overline{k}(\mathbf{s}^{\prime},\tau^{\prime})$ ensures that $\frac{k^{\prime}}{\overline{k}(\mathbf{s}^{\prime},\tau^{\prime})}G_{\mathbf{s}^{\prime}}(k^{\prime},\tau^{\prime}|0,0)$ is properly normalized. Equation (6) implies that $k$ and $k^{\prime}$ are conditionally independent given the state and age, with their overall correlation mediated by the marginal dependencies in the joint state-age distribution. In other words, degree correlation is entirely encoded in the state-age correlation within $\mathbf{\Psi}$.
@@
Empirical Validation. To validate our theory against empirical data, we applied it to the scientific innovation system, which can be understood as a causal network, where the creation of innovations causally depends on existing knowledge. We used the Web of Science (WOS) dataset spanning 1970–2014, comprising 44M publications and 800M citations. For simplicity, the causal network considered here is the citation network, focusing on events of paper publication with a single type of causal relationship $\mathbf{s}_{\text{obs}}=$ “cite”, which links new knowledge to its intellectual predecessors. Using the WOS dataset, we extracted four citation networks across different disciplines, including biology, chemistry, mathematics, and physics, comprising 13M, 9M, 3M, and 7M vertices, respectively (see SM Section 2.1).
@@
Figure 1: Empirical results (scatter points) vs. theoretical predictions (solid lines) for (a) The network size $n(t)$ of citation graphs across biology, chemistry, mathematics, and physics. Theoretical predictions are shown as straight lines with predicted growth rates. (b) The probability distribution of states $\psi(\mathbf{s})$. (c) The degree distribution $P(k)$. (d) The empirically measured causal correlation kernel $K(\lambda|\lambda^{\prime},\tau)$ for physics citation graphs, rescaled by its average value $\overline{\lambda}(\lambda^{\prime},\tau)$. The rescaled kernel collapses onto a master curve. Similar plots for other subjects are provided in the Supplementary Material.
@@
Figure 2: Normalized joint degree distribution for (a–b) biology, (c–d) chemistry, (e–f) mathematics, and (g–h) physics. The left and right columns represent empirical measurements and theoretical predictions, respectively.
Existing studies suggest that the growth process at the individual level follows a reinforced Poisson process (RPP), incorporating both preferential attachment and aging [41, 12]. Specifically, the growth rate is given by $\Lambda{(k,\tau)}=\lambda\phi(\tau)(k+k_{0})$, where $\lambda$ represents fitness. The aging function, typically modeled as a log-normal distribution $\phi(t)=\frac{1}{\sqrt{2\pi}\sigma t}\exp\left[-\frac{(\ln{t}-\mu)^{2}}{2\sigma^{2}}\right]$, captures the long-tail dynamic correlation. The corresponding Green’s function is given by the negative binomial distribution (see SM Section 1.1), with the average growth
$\overline{k}(\lambda,t)=k_{0}\left(\exp\left(\lambda\int_{0}^{t}\phi(t^{\prime})\mathrm{d}t^{\prime}\right)-1\right),$ (7)
which implies a finite ultimate degree $\overline{k}_{\infty}(\lambda)\mathrel{\mathop{:}}=\overline{k}(\lambda,\infty)=k_{0}\left(e^{\lambda}-1\right)$. While Eq. (7) successfully captures individual citation growth, it has several drawbacks: (1) it does not account for causal correlations, and (2) it assigns the parameters $\lambda$, $k_{0}$, $\mu$, and $\sigma$ to each paper, leading to a large number of free parameters that are experimentally determined. In particular, the number of parameters scales with the number of vertices, $O(N)$. Such limitations are well-documented in fitness models [24, 25].
@@
Our theory addresses the above drawbacks by treating individual parameters as latent states, $\mathbf{s}_{\text{lat}}$. This approach allows our theory (4) to predict the distribution of these latent states, thereby avoiding the large parameter space. To solve Eq. (4), it is necessary to determine the causal correlation kernel, $K$. For simplicity, we focus solely on the fitness parameter $\mathbf{s}_{\text{lat}}=\lambda$, keeping all other parameters $\sigma=1$ and $\mu=2.3$ fixed globally. Figure 1d plots a rescaled $K(\lambda|\lambda^{\prime},t)$, finding that it collapses into master curves after scaling by its mean $\overline{\lambda}$, as
$K(\lambda|\lambda^{\prime},t)=\frac{1}{\overline{\lambda}(\lambda^{\prime},t)}\mathcal{K}\left(\frac{\lambda}{\overline{\lambda}(\lambda^{\prime},t)}\right).$ (8)
Here, the mean fitness $\overline{\lambda}(\lambda^{\prime},t)$ is an exponential function of the ratio $\overline{k}(\lambda^{\prime},t)/\overline{k}_{\infty}(\lambda^{\prime})$, as shown in our previous work [42]. Moreover, the universal distribution $\mathcal{K}$ is best fitted by the mixed Weibull distribution function (see SM Section 2.3).
@@
Substituting Eqs. (78) into Eq. (4), our theory predicts the network growth rates $r\approx 0.037$, $0.03$, $0.04$, and $0.028$ for biology, chemistry, mathematics, and physics, respectively, which are in excellent agreement with the empirical measurements (Fig. 1a). Figure 1b compares the predicted fitness distribution $\psi(\lambda)$ with the empirical data, showing remarkable consistency. Here, the empirical $\psi(\lambda)$ is obtained by fitting the RPP model to individual papers. Additionally, the degree distribution $P(k)$ is computed analytically using Eq. (5), again aligning perfectly with experimental results (Fig. 1c).
To validate the model’s prediction of topical correlation, we calculate the joint degree distribution $P(k,k^{\prime})$, using Eq. (6). Figure 2 plots the normalized joint degree distribution $R(k,k^{\prime})\mathrel{\mathop{:}}=\frac{P(k,k^{\prime})}{P_{rand}(k,k^{\prime})}$, i.e., $P(k,k^{\prime})$ compared to the randomized one [28]. Our results show that the theoretical prediction successfully captures the assortative nature of the citation graph, aligning well with empirical measurements.
Figure 3: K-nearest-neighbor $k_{nn}(k^{\prime})$ for both empirical measurements (scatter points) and theoretical predictions (solid curves) across four subjects.
To compare our theory with empirical data more quantitatively, we further calculate the $k$-nearest-neighbor function, defined as $k_{nn}(k^{\prime})\mathrel{\mathop{:}}=\sum_{k}kP(k|k^{\prime})$, which measures the dependency of a vertex’s degree on that of its neighbors. Based on Eq. (6), we obtain (see SM Section 1.4)